# Measuring software innovation with open source software development data

### v1

@@ Abstract
This paper introduces a novel measure of software innovation based on open source software (OSS) development activity on GitHub. We examine the dependency growth and release complexity among $\sim$200,000 unique releases from 28,000 unique packages across the JavaScript, Python, and Ruby ecosystems over two years post-release. We find that major versions show differential, strong prediction of one-year lagged log change in dependencies. In addition, semantic versioning of OSS releases is correlated with their complexity and predict downstream adoption. We conclude that major releases of OSS packages count as a unit of innovation complementary to scientific publications, patents, and standards, offering applications for policymakers, managers, and researchers.
@@ 1 Introduction
There is widespread, interdisciplinary research and policy interest in measuring innovation, with attention traditionally focusing on academic outputs [1], business outcomes [2], and their interactions [3]. Yet software innovation, particularly in open source software (OSS) innovation, often eludes conventional metrics like scientific publications and patents. While OSS are digital public goods that firms leverage as an “inward knowledge flow” to produce innovations [2], it also constitutes innovation in its own right. The OECD Oslo Manual [2] frames innovation through product and process advances, evaluating novelty either by comparing to the current state of the art or by assessing ex-post impact. This framework poses unique challenges for OSS, where individual contributions may appear incremental—often just a few lines of code—but collectively drive radical innovation and creative destruction [4]. This tension between granular contributions and systemic impact highlights the need for new approaches to measuring OSS innovation.
@@
Previously, researchers have sought to measure OSS innovation through collaboration activity [5, 6], its economic impact [7, 8], and value [9, 10]. In this paper, we define and operationalize a novel method of measuring OSS innovation in discrete units, offering researchers and policymakers a software innovation metric that is complementary to publications, patents, and standards. Our method uses publicly available data from GitHub, leveraging software developers’ publication of packages and others developers’ decisions to adopt the software. Developers often use semantic versioning when publishing a change to a software package that denotes its significance as major, minor, or a patch [11]. In practice, then, versioning increments may represent inventive steps in software. By reviewing other developers’ activity in response to new versions, the open source community can offer external validation of whether an increment is in fact innovative. Specifically, we assess package dependencies by examining how frequently new versions are integrated into other projects over time, providing a metric to evaluate the impact of these inventive steps.
@@
Data & methods: We analyze a dataset of approximately 200,000 unique releases from 28,000 unique packages across the JavaScript, Python, and Ruby ecosystems on GitHub, examining package dependencies and release complexity over time frames up to two years post-release. Our methodology combines quantitative analysis of dependency growth patterns with a machine learning approach to assess release complexity using GPT-4.
Key finding 1: At the one-year post-release mark, major releases are associated with significantly larger increases in package dependencies compared to minor or patch releases across the JavaScript, Python, and Ruby ecosystems, highlighting distinct patterns of dependency growth that emerge following significant package updates.
Key finding 2: Major releases continue to generate larger positive changes in log-difference compared to minor and patch releases, regardless of the package ecosystem, initial package utilization, or time frame examined up to two years post-release. However, the magnitude of the log-difference tends to diminish slightly in the second year.
Key finding 3: Release complexity, as rated by a prompted GPT-4 model, correlates slightly positively with dependency growth, particularly for JavaScript and Python packages. However, the Ruby ecosystem does not exhibit the same correlation, suggesting different dynamics in the Ruby community and a different handling of release notes.
Limitations: We do not account for software licensing effects on dependency relationships (e.g. GPL vs. Apache v2), geographic patterns of adoption, or the inherent variability in our LLM-based semantic analysis methodology.
Implications statement: The findings provide new insights into OSS innovation dynamics, offering potential applications for policymakers, practitioners, and researchers in evaluating and fostering innovation in OSS development.
@@
The paper is structured as follows. In § 2, we discuss prior work on innovation indicators and the gap for OSS innovation. In § 3, we discuss our approach to the measurement of units of software innovation using GitHub data. In § 3, we report our key findings. In § 4, we discuss the implications and limitations of our method, as well as future research directions.
@@ 2.1 Innovation indicators and the open source gap
Additionally, underpinning indicators include methodologies for measuring innovation, such as the selection of units of innovation, with novelty being a key consideration. The OECD defines novelty as a criterion for determining whether a product or process is “significantly different” [2] from previous iterations, thus qualifying as an innovation. This evaluation can be approached in two ways: first, by comparing a firm’s innovations with the state of the art within the firm and its market or industry; and second, by assessing the impact of an innovation in transforming or creating a new market, which may indicate radical or disruptive innovation. Unlike the first approach, assessing the impact of an innovation can only be done ex-post, often requiring several years to evaluate its effects.
Hence, the 2018 version of the OECD Oslo Manual advocates for the inclusion of novel digital data sources, emphasizing how automated data collection can expand the range of available data, thereby enhancing the development of more diverse and dynamic innovation indicators [2] [20]. Traditional indicators, in particular, often lack timeliness and accuracy and tend to underestimate innovation activities in some contexts [20]. In response, recent studies have begun to explore the potential of novel indicators, for example, based on web-based data sources such as startup databases and online repositories (see e.g. [21] [20]).
@@
Recently, research interest in the economics of OSS has begun to increase, albeit not for OSS innovation indicators directly. Hoffman et al. (2024) [9] estimate the supply-side value of OSS at \$4.15 billion, while the demand-side value is estimated to reach a significantly larger \$8.8 trillion. The authors conclude that firms would need to spend 3.5 times more on software if OSS did not exist, highlighting not only its economic impact of OSS but also its innovative potential in software development. Meanwhile Korkmaz et al. (2024) [10] develop a framework for measuring OSS value using GitHub data, estimating US investment in OSS at \$37.8 billion in 2019. Furthermore, Wright et al. (2023) [8] demonstrate a correlation between country-level GitHub participation and increased founding of innovative ventures, suggesting that OSS activity serves as a catalyst for innovation-driven entrepreneurship. Lin and Maruping (2021) [23] add nuance to this relationship, showing that the benefits of OSS engagement for startups vary across their lifecycle stages, with different advantages accruing from inbound and outbound OSS activities.
@@ 2.2 Learning from existing innovation indicators
In this section, we establish connections between established innovation indicators and potential OSS innovation indicators. Scientific publications offer a well-established benchmark for innovation measurement. Key performance indicators include the volume of output and forward citations received [36]. Brown et al. demonstrated a similar approach for measuring the impact of scientific OSS repositories and software projects by creating large networks of scientific publications with links to software used and mentioned in each article [37]. This approach aligns with traditional bibliometric analyses but applies them to software artifacts.
Patents provide another crucial external indicator. They can reference OSS projects, particularly in non-patent literature (NPL) citations, indicating the industrial relevance and potential commercial value of a repository. As Harhoff et al. (2003) [38] note, NPL citations are particularly important in fields like pharmacy and chemistry, while in the ICT sector standards are also frequently cited [39]. For OSS repositories, one can track the number of patents citing them and analyze the characteristics of the citing patents and their applicants.
@@ 2.3 Prior approaches to measuring OSS innovation
GitHub offers rich data for OSS innovation measurement. The platform provides various indicators for the performance and popularity of OSS repositories, including commits, stars [62], watching [63], forks [64], and pull requests [65]. Project metrics, in particular repository forks, have been shown to correlate with usage [66], offering a proxy variable for scarce usage and adoption data. The GitHub dependency graphs illustrates the projects a repository depends on and those that depend on it, offering a network view of a project’s influence and reliance in the broader software ecosystem [56]. Table LABEL:tab:oss-indicator-characteristics discusses the merits and limitations of these internal indicators, and Table LABEL:tab:comparison-patents-standards-oss shows key variables for indicators from OSS repositories, patents, and standards.
@@ 3.2.1 Data Generation
We constructed our dataset from internal GitHub databases. Most of the data, such as repository stars, forks, and programming languages, can be accessed via the GitHub API. Additionally, dependent data can be sourced from third-party packages, such as https://github.com/nvuillam/github-dependents-info or from services such as https://ecosyste.ms/. However, results from these third-party options may vary in dependent counts due to differing counting criteria and the absence of look-ahead or look-behind metrics, which are essential for comprehensive analysis reproduction. While public sources support future use of our method, to our knowledge, no third-party has attempted a comprehensive collection of daily snapshots of repository metadata and dependency graph information for the time period considered in our study. Our dataset, available at https://github.com/evamaxfield/dg-uoi-analysis, is the combination of three datasets of innovation-related activities on the GitHub platform.
@@
The second dataset focuses on package releases. This data provides key information about the versions of packages that are being updated, with each entry including a repository identifier alongside release notes describing changes for each version. For instance, an entry includes the account name, the repository name, and the release version of a Python package with a specific name, along with the corresponding date. For example, ”my-account/my-repo” refers to the account and repository name, while ”v1.2.3” indicates the package version, and the release date is September 25, 2024). This information is detected and logged daily through GitHub’s automatic parsing of standard packaging files, such as pyproject.tom and setup.py for Python packages, and package.json for JavaScript packages.
@@
The third dataset focuses on dependents of a package, which identifies the down-stream repositories that rely on specific packages along with their corresponding dependency versions. Each entry in this dataset includes a repository identifier for the dependent repository, the name of the package it depends on, and the version of that dependency. For instance, an entry indicates that ”another-account/another-repo” depends on a Python package named ”zyx,” with the version ”v1.2.2,” as of September 25, 2024. Similarly, this information is detected and logged each day via GitHub’s automatic parsing of standard packaging files (see above).
@@
We combine and filter these datasets to create a comprehensive view of package releases and their dependencies. First, for each day of newly detected package releases, we join each package release with the corresponding repository information. Next, we implement a filtering mechanism to enhance data quality. Specifically, we remove package releases originating from repositories that do not have at least one star or are classified as “forked repositories”. This step is crucial to eliminate packages that may be either extremely new, lacking community support, have not undergone sufficient development, or may be low-quality or automatically generated with little meaningful content (e.g. spam). This ensures that our dataset reflects more established and relevant packages, focusing on those with active development and community engagement. We apply another filter to exclude package releases that do not adhere to the semantic versioning specification (see § 3.1 above). During this process, we decompose the version string and categorize releases accordingly to major, minor, and patch. This classification allows us to systematically analyze the frequency and impact of updates based on the significance of the changes.
@@
Next, we filter to ensure that only package releases from repositories whose names align with their respective package names are included in our dataset. Identifying the primary package for a repository is crucial because it simplifies the understanding of package dependencies and updates. For instance, the repository https://github.com/pandas-dev/pandas contains the source code for the widely used Python package “pandas,” and thus, its package releases are included in our dataset. Conversely, if the repository were named “pandas-dev/pandas-subpkgs,” where the repository name (“pandas-subpkgs”) does not match the package name (“pandas”), those releases would be excluded. This filtering criterion helps avoid confusion that may arise when multiple packages are housed within a single repository, making it difficult to track dependencies accurately and understand which package is being referred to in analyses. By applying this criterion, we aim to focus on the packages that are directly associated with their respective repositories, thereby enhancing the quality and reliability of our dataset.
@@
Lastly, we enrich our dataset by calculating several measures. First we integrate look-ahead metric information regarding repository stars and forks. This involves systematically retrieving and compiling repository data for each specified look-ahead day interval per repository. Specifically, for each repository, we collect data on the number of stars and forks over the defined intervals, which serves as an indicator of community engagement and interest. Second, we integrate the number of package dependents. This is accomplished by counting the instances of non-forked repositories that have received at least one star and include a dependency on a package sharing the same name, within the same ecosystem (e.g. Python, JavaScript), and aligned with the same major version number.
@@ 3.2.2 Data Processing
Our dataset, compiled from dependency graph package release detection (see § 3.2.3, in total includes 862,653 software releases from 148,628 unique packages from multiple different software ecosystems on GitHub. Among these, the most common ecosystems are JavaScript (NPM), Python (PyPI), and Ruby (RubyGems).
Following dataset compilation, we implemented a series of necessary data processing  [^1] All of these pre-processing activities were executed using the Polars library in Python [67], which facilitated efficient manipulation of large datasets and streamlined the data processing workflow steps to prepare the data for analysis. First, we removed any packages that had multiple releases on the same day. While it is common for packages to have multiple releases over longer time horizons, excluding same-day versions minimizes the risk of applying the same set of metrics to different releases. [^2] This scenario typically arises in two contexts: (i) a new major or minor release of a package which may be followed by a patch release to rectify bugs inadvertently introduced with new features, and (ii) mature software projects may maintain multiple versions of the same software, such as a ”Long-Term Support” (LTS) version alongside a leading-edge version. Second, we excluded packages not belonging to the JavaScript, Python, and Ruby ecosystems, as these languages consistently accounted for at least 5,000 package releases across all datasets (spanning six-month, one-year, and two-year intervals). Third, we also excluded any packages that did not have at least one dependent prior to the release. This criterion ensures that only actively utilized packages are included in our analysis, enhancing the focus on software that has demonstrated relevance within the ecosystem.
Following these steps, our filtered dataset 202,014 unique releases from 28,044 unique packages from the JavaScript (NPM), Python (PyPI), and Ruby (RubyGems) ecosystems. [^3] Release and package counts are based on the produced one-year look-ahead metric dataset. Different look-ahead time-span datasets resulted in different counts however we focused on the final counts for the one-year look-ahead metrics because our analysis also focuses on the same portion of the data.
Subsequently, we stratified the dataset based on package popularity, measured by existing dependents. We categorized the packages into three bins: those in the less than 25th percentile, those between the 25th and 75th percentiles, and those in the greater than 75th percentile of package dependents prior to release. Additionally, we classified the packages by release series based on their versioning: zero-version, one-version, and two-plus-version (i.e., 0., 1., and 2+.*). This dual stratification reflects both the popularity and maturity of the packages.

### v2

@@ Abstract
Existing innovation metrics inadequately capture software innovation, creating blind spots for researchers and policymakers seeking to understand and foster technological innovation in an increasingly software-defined economy. This paper introduces a novel measure of software innovation based on open source software (OSS) development activity on GitHub. We examine the dependency growth and release complexity among $\sim$350,000 unique releases from 33,000 unique packages across the JavaScript, Python, and Ruby ecosystems over two years post-release. We find that the semantic versioning types of OSS releases exhibit ecosystem-specific and maturity-dependent patterns in predicting one-year dependency growth, with minor releases showing relatively consistent adoption across contexts while major and patch releases vary significantly by ecosystem and package size. In addition, while semantic versioning correlates with the technical complexity of the change-set, complexity itself shows minimal correlation with downstream adoption, suggesting that versioning signals rather than technical change drive dependency growth. Overall, while semantic versioning release information can be used as a unit of innovation in OSS development complementary to common sources for innovation metrics (e.g. scientific publications, patents, and standards), this measure should be weighted by ecosystem culture, package maturity, and release type to accurately capture innovation dynamics. We conclude with a discussion of the theoretical and practical implications of this novel measure of software innovation as well as future research directions.
@@ 1 Introduction
There is interdisciplinary research and policy interest in measuring innovation, with attention traditionally focusing on academic outputs [1], business outcomes [2], and their interactions [3]. Yet, software innovation, particularly open source software (OSS) innovation, often eludes conventional metrics such as scientific publications, patents, and standards. Although OSS are digital public goods that firms leverage as an “inward knowledge flow” to produce innovations [2], it also constitutes innovation in its own right. The OECD Oslo Manual [2] frames innovation through product and process advances, evaluating novelty either by comparing to the current state of the art or by assessing ex-post impact. This framework poses unique challenges for OSS, where individual contributions may appear incremental—often just a few lines of code—but collectively drive radical innovation and creative destruction [4]. This tension between granular contributions and systemic impact highlights the need for new approaches to measuring OSS innovation.
@@
Previously, researchers have sought to measure OSS innovation through collaboration activity [5, 6], its economic impact [7, 8], and economic value [9, 10]. In this paper, we define and operationalize a novel method for measuring OSS innovation in discrete units, offering researchers and policymakers a software innovation metric that is complementary to traditional sources for innovation metrics. Our method uses publicly available data from GitHub, leveraging software developers’ publication of packages and other developers’ decisions to adopt the software. Developers often use semantic versioning when publishing a change to a software package that denotes its significance as major, minor, or patch [11]. In practice, then, versioning increments may represent inventive steps in software. By reviewing other developers’ activity in response to new versions, the open source community can offer external validation of whether an increment is, in fact, innovative. Specifically, we assess package dependencies by examining how frequently new versions are integrated into other projects over time, providing a metric to evaluate the impact of these inventive steps.
Problem statement: Conventional innovation metrics inadequately capture the granular, iterative, yet collectively transformative nature of software innovation, creating blind spots for researchers, software managers, and policymakers seeking to understand and support technological innovation in an increasingly software-driven economy.
@@
Data & methods: We analyze a dataset of approximately 350,000 unique releases from 33,000 unique packages across the JavaScript, Python, and Ruby ecosystems on GitHub, examining package dependencies and release complexity over time frames up to two years post-release. Our methodology combines quantitative analysis of dependency growth patterns with a machine learning approach to assess release complexity using GPT-4.
Key finding 1: Semantic versioning types exhibit ecosystem-specific and pre-release utilization-dependent patterns in predicting dependency growth at both one-year and two-year post-release time frames. Minor releases demonstrate relatively consistent adoption benefits across JavaScript, Python, and Ruby ecosystems and initial dependency levels. In contrast, major and patch releases vary significantly based on ecosystem culture and baseline package popularity.
Key finding 2: Ecosystem-specific adoption patterns remain consistent throughout all stages of a package’s development, from early (pre-1.0) to mature (2.0 and beyond). Across these stages, minor releases reliably achieve steady adoption, highlighting their role as innovation signals for OSS.
Key finding 3: A release’s technical complexity correlates with semantic versioning type but shows minimal direct correlation with dependency growth, indicating that versioning serves as a signaling mechanism that drives adoption independently of actual technical complexity.
Methodological implications: These ecosystem-specific and context-dependent patterns suggest that future innovation measurement should normalize adoption metrics by ecosystem culture, baseline package popularity, and development lifecycle stage, rather than treating semantic versioning uniformly across contexts. Furthermore, by operationalizing dependency growth following semantic versioned releases as a scalable, real-time proxy for software innovation, our approach provides researchers, practitioners, and policymakers with a transparent and reproducible metric that complements traditional innovation indicators and enables more granular analysis of digital innovation dynamics .
Limitations: We do not account for software licensing effects on dependency relationships (e.g. GPL vs. Apache v2), or geographic patterns of adoption. These aspects were excluded due to limitations in the available data and the complexity of reliably linking licenses and geographic meta data to dependency events at scale. Additionally, we did not address the the inherent variability in our LLM-based semantic analysis methodology, as we prioritized scalability and automation over manual validation, considering the large data set. Future research could address these dimensions to further refine and contextualize software innovation metrics
Impact statement: Our findings provide new insights into OSS innovation dynamics, by revealing ecosystem-sensitive metrics that capture the adoption and influence of OSS releases. This enables researchers to more accurately measure digital innovation, equips practitioners with data-driven tools for strategic decision-making, and empowers policymakers to design more effective interventions for fostering OSS-driven technological progress.
@@
The paper is structured as follows. In § 2, we discuss prior work on innovation indicators and the gap for OSS innovation. In § 3, we discuss our approach to measuring units of software innovation using GitHub data. In § 4, we report our key findings. In § 5, we discuss the implications and limitations of our method, as well as future research directions.
@@ 2.1 Innovation indicators and the open source gap
Additionally, underpinning indicators include methodologies for measuring innovation, such as the selection of units of innovation, with novelty being a key consideration. The OECD defines novelty as a criterion for determining whether a product or process is “significantly different” [2] from previous iterations, thus qualifying as an innovation. This evaluation can be approached in two ways: first, by comparing a firm’s innovations with the state-of-the-art within the firm and its market or industry; and second, by assessing the impact of an innovation in transforming or creating a new market, which may indicate radical or disruptive innovation. Unlike the first approach, assessing the impact of an innovation can only be done ex-post, often requiring several years to evaluate its effects.
Hence, the 2018 version of the OECD Oslo Manual advocates for the inclusion of novel digital data sources, emphasizing how automated data collection can expand the range of available data, thereby enhancing the development of more diverse and dynamic innovation indicators [2] [20]. Traditional indicators, in particular, often lack timeliness and accuracy and tend to underestimate innovation activities in some contexts [20]. In response, recent studies have begun to explore the potential of novel indicators, for example, based on web-based data sources such as startup databases and online repositories (see e.g. [21] [22] [20]).
@@
Recently, research interest in the economics of OSS has begun to increase, albeit not for OSS innovation indicators directly. Hoffmann et al. (2024) [9] estimate the supply-side value of OSS at \$4.15 billion, while the demand-side value is estimated to reach a significantly larger \$8.8 trillion. The authors conclude that firms would need to spend 3.5 times more on software if OSS did not exist, highlighting not only its economic impact of OSS but also its innovative potential in software development. Meanwhile, Korkmaz et al. (2024) [10] develop a framework for measuring OSS value using GitHub data, estimating US investment in OSS at \$37.8 billion in 2019. Furthermore, Wright et al. (2023) [8] demonstrate a correlation between country-level GitHub participation and increased founding of innovative ventures, suggesting that OSS activity serves as a catalyst for innovation-driven entrepreneurship. Lin and Maruping (2021) [23] add nuance to this relationship, showing that the benefits of OSS engagement for startups vary across their lifecycle stages, with different advantages accruing from inbound and outbound OSS activities.
@@ 2.2 Learning from existing innovation indicators
In this section, we establish connections between established innovation indicators and potential OSS innovation indicators. Scientific publications offer a well-established benchmark for innovation measurement. Key performance indicators include the volume of output and forward citations received [36]. Brown et al. (2023) demonstrated a similar approach for measuring the impact of scientific OSS repositories and software projects by creating large networks of scientific publications with links to software used and mentioned in each article [37]. This approach aligns with traditional bibliometric analyses but applies them to software artifacts.
Patents provide another crucial external indicator. They can reference OSS projects, particularly in non-patent literature (NPL) citations, indicating the industrial relevance and potential commercial value of a repository. As Harhoff et al. (2003) [38] note, NPL citations are particularly important in fields such as pharmacy and chemistry, while in the ICT sector standards are also frequently cited [39]. For OSS repositories, one can track the number of patents citing them and analyze the characteristics of the citing patents and their applicants.
@@ 2.3 Prior approaches to measuring OSS innovation
GitHub offers rich data for OSS innovation measurement. The platform provides various indicators for the performance and popularity of OSS repositories, including commits, stars [62], watching [63], forks [64], and pull requests [65]. Project metrics, in particular repository forks, have been shown to correlate with usage [66], offering a proxy variable for usage and adoption data. The GitHub dependency graphs illustrates the projects a repository depends on and those that depend on it, offering a network view of a project’s influence and reliance in the broader software ecosystem [56]. Table LABEL:tab:oss-indicator-characteristics discusses the merits and limitations of these internal indicators, and Table LABEL:tab:comparison-patents-standards-oss shows key variables for indicators from OSS repositories, patents, and standards.
@@ 3.2.1 Data Generation
We constructed our dataset from internal GitHub databases. Most of the data, such as repository stars, forks, and programming languages, can be accessed through the GitHub API. Additionally, dependent data can be sourced from third-party packages, such as https://github.com/nvuillam/github-dependents-info or from services such as https://ecosyste.ms/. However, results from these third-party options may vary in dependent counts due to differing counting criteria and the absence of look-ahead or look-behind metrics, which are essential for comprehensive analysis reproduction. While public sources support future use of our method, to our knowledge, no third party has attempted a comprehensive collection of daily snapshots of repository metadata and dependency graph information for the time period considered in our study. Our dataset, available at https://github.com/evamaxfield/dg-uoi-analysis, is the combination of three datasets of innovation-related activities on the GitHub platform.
@@
The second dataset focuses on package releases. This data provides key information about the updated versions of packages, with each entry including a repository identifier alongside release notes describing changes for each version. For instance, an entry includes the account name, the repository name, the release version of a Python package with a specific name, and the corresponding date. For example, ”my-account/my-repo” refers to the account and repository name, while ”v1.2.3” indicates the package version, and the release date is September 25, 2024. This information is detected and logged daily through GitHub’s automatic parsing of standard packaging files, such as pyproject.tom and setup.py for Python packages, and package.json for JavaScript packages.
@@
The third dataset focuses on dependents of a package, which identifies the downstream repositories that rely on specific packages. Each entry in this dataset includes a repository identifier for the dependent repository and the name of the package on which it depends. For instance, an entry indicates that ”another-account/another-repo” depends on a Python package named ”zyx” as of September 25, 2024. Similarly, this information is detected and logged daily via GitHub’s automatic parsing of standard packaging files (see above).
@@
We combine and filter these datasets to create a comprehensive view of package releases and their dependencies. First, we join each package release with the corresponding repository information for each day of newly detected package releases. Next, we implement a filtering mechanism to enhance data quality. Specifically, we remove package releases originating from repositories that do not have at least one star or are classified as “forked repositories”. This step is crucial to eliminate packages that may be either extremely new, lacking community support, have not undergone sufficient development, or may be low-quality or automatically generated with little meaningful content (e.g. spam). This ensures that our dataset reflects more established and relevant packages, focusing on those with active development and community engagement. We apply another filter to exclude package releases that do not adhere to the semantic versioning specification (see § 3.1 above). During this process, we decompose the version string and categorize releases accordingly to major, minor, and patch. This classification allows us to systematically analyze the frequency and impact of updates based on the significance of the changes.
@@
Next, we filter to ensure that only package releases from repositories whose names align with their respective package names are included in our dataset. Identifying the primary package for a repository is crucial because it simplifies understanding package dependencies and updates. For instance, the repository https://github.com/pandas-dev/pandas contains the source code for the widely used Python package “pandas,” thus, its package releases are included in our dataset. Conversely, if the repository were named “pandas-dev/pandas-subpkgs,” where the repository name (“pandas-subpkgs”) does not match the package name (“pandas”), those releases would be excluded. This filtering criterion helps avoid confusion when multiple packages are housed within a single repository, making it challenging to track dependencies accurately and understand which package is being referred to in analyses. By applying this criterion, we aim to focus on the packages directly associated with their respective repositories, thereby enhancing the quality and reliability of our dataset.
@@
Lastly, we enrich our dataset by calculating several measures. First, we integrate look-ahead metrics regarding repository stars and forks. This involves systematically retrieving and compiling repository data for each specified look-ahead day interval per repository. Specifically, for each repository, we collect data on the number of stars and forks over the defined intervals, which serves as an indicator of community engagement and interest. Second, we integrate the number of package dependents. This is accomplished by counting the instances of non-forked repositories that have received at least one star and include a dependency on a package sharing the same name, and within the same ecosystem (e.g. Python, JavaScript).
@@ 3.2.2 Data Processing
Our dataset, compiled from dependency graph package release detection (see § 3.2.3, in total includes 1,492,270 software releases from 167,375 unique packages from multiple different software ecosystems on GitHub. Among these, the most common ecosystems are JavaScript (NPM), Python (PyPI), and Ruby (RubyGems).
Following dataset compilation, we implemented a series of necessary data processing  [^1] All of these pre-processing activities were executed using the Polars library in Python [67], which facilitated efficient manipulation of large datasets and streamlined the data processing workflow steps to prepare the data for analysis. First, we removed any packages with multiple releases on the same day. While it is common for packages to have multiple releases over longer time horizons, excluding same-day versions minimizes the risk of applying the same set of metrics to different releases. [^2] This scenario typically arises in two contexts: (i) a new major or minor release of a package which may be followed by a patch release to rectify bugs inadvertently introduced with new features, and (ii) mature software projects may maintain multiple versions of the same software, such as a ”Long-Term Support” (LTS) version alongside a leading-edge version. Second, we excluded packages not belonging to the JavaScript, Python, and Ruby ecosystems, as these languages consistently accounted for at least 5,000 package releases across all datasets (spanning six-month, one-year, and two-year intervals). Third, we excluded any packages that did not have at least five dependents before the release. This criterion ensures that only actively utilized packages are included in our analysis, enhancing the focus on software that has demonstrated relevance within the ecosystem.
Following these steps, our filtered dataset consists of 356,220 unique releases from 33,236 unique packages from the JavaScript (NPM), Python (PyPI), and Ruby (RubyGems) ecosystems. [^3] Release and package counts are based on the produced one-year look-ahead metric dataset. Different look-ahead time-span datasets resulted in different counts; however, we focused on the final counts for the one-year look-ahead metrics because our analysis also focuses on the same portion of the data.
Subsequently, we stratified the dataset based on package popularity, measured by existing dependents. We categorized the packages into four log-scale bins: packages with less than 100 dependents (small), packages with between 100 and 1,000 dependents (medium), packages with between 1,000 and 10,000 dependents (large), and packages with greater than 10,000 dependents (huge). Additionally, we classified the packages by release series based on their versioning: zero-version (i.e., 0.*), one-version (i.e., 1.*), and two-plus-version (i.e., 2+.*). This dual stratification reflects both the popularity and maturity of the packages.